The formula h=-16t^2 + v0t +s0 give sthe hieght of an object tossed upward where v0 represents the initial velocity, s0 represents the initial height, and t represents time. A golf ball is his straight up from the ground level with an initial velocity of 72 ft/sec.

Found 2 solutions by ikleyn, johnsss:
Answer by ikleyn(52781)

.
The formula h=-16t^2 + v0t +s0 give the height of an object tossed upward where v0 represents
the initial velocity, s0 represents the initial height, and t represents time.
A golf ball is his straight up from the ground level with an initial velocity of 72 ft/sec.
~~~~~~~~~~~~~~~~~~~~~~~~


You forgot to attach your question at the end of your post.

My warm parting words to you are "read what you post before you submit your message".

Be accurate, respect the tutors and respect yourself.



Answer by johnsss(2)

To solve the problem using the equation.
Source: https://mipromedio.co/blog/como-calcular-promedio-universitario/

=

16
𝑡
2
+
𝑣
0
𝑡
+
𝑠
0
h=−16t
2
+v
0

t+s
0

, let's break it down:
Given:
𝑣
0
=
72
v
0

=72 ft/sec (initial velocity)
𝑠
0
=
0
s
0

=0 (since the ball is hit from ground level)
The equation becomes

=

16
𝑡
2
+
72
𝑡
h=−16t
2
+72t
This formula describes the height of the golf ball at any given time
𝑡
t. If you need to find specific values, like the maximum height or when the ball hits the ground, you'd follow these steps:
Maximum Height: The maximum height occurs at the vertex of the parabola, which can be found by the formula
𝑡
max
=

𝑣
0
2
𝑎
t
max

=−
2a
v
0



. For this equation,
𝑎
=

16
a=−16 and
𝑣
0
=
72
v
0

=72, so:
𝑡
max
=

72
2
(

16
)
=
72
32
=
2.25

seconds
t
max

=
2(−16)
−72

=
32
72

=2.25seconds
Then substitute this into the equation to find the height at this time:

=

16
(
2.25
)
2
+
72
(
2.25
)
=

16
(
5.0625
)
+
162
=

81
+
162
=
81

ft
h=−16(2.25)
2
+72(2.25)=−16(5.0625)+162=−81+162=81ft
So, the maximum height is 81 feet.
Time to hit the ground: The ball will hit the ground when

=
0
h=0. Set the height equation equal to zero and solve for
𝑡
t:
0
=

16
𝑡
2
+
72
𝑡
0=−16t
2
+72t
Factor:
0
=
𝑡
(

16
𝑡
+
72
)
0=t(−16t+72)
So,
𝑡
=
0
t=0 (at launch) or
𝑡
=
72
16
=
4.5

seconds
t=
16
72

=4.5seconds.